How much heat is lost when 350g of iron cools from 54.1 to 22.4.
1
Expert's answer
2018-03-28T05:26:28-0400
m = 0.35 kg
C(Fe) = 450 J/(kg*K)
T1=54.1oC
T2=22.4oC
Q = C*m*(T1-T2)
Q= 450*0.35*31.7 = 4992.75 (J)